OpenStudy (saylilbaby):

The perimeter of a rectangle is 50 feet and its area is 154 feet. What are the dimensions of the rectangle?

Vocaloid (vocaloid):

Well, we know that P = 2(l+w) and A = lw, right? with me so far?

Vocaloid (vocaloid):

A = lw = 154, so we can solve for L by dividing both sides by W

Vocaloid (vocaloid):

L = 154/w, then we can plug this into the perimeter equation and solve for w

Vocaloid (vocaloid):

hold on a sec, lemme check your math

Vocaloid (vocaloid):

hm, not quite, can you show me all your work?

Vocaloid (vocaloid):

ok, so if L = 154/W, and P = 2(L+W), we substitute in "154/W" for L, can you tell me what you get after you do that?

OpenStudy (saylilbaby):

@Abhisar @Michele_Laino